OpenStudy (anonymous):

sin 2x=.85 (Using a graphing calculator, find solutions 0 ≤ x ≤2π and express solutions accurate to 3 sig. figures)

OpenStudy (anonymous):

I already have it in radian mode but the answer I'm getting is different from the one I'm seeing in the back of my book where all the answers are.

OpenStudy (anonymous):

x=29.105 degree

OpenStudy (anonymous):

there's suppose to be a couple of solutions...

OpenStudy (anonymous):

second is 180-29.105

OpenStudy (anonymous):

oh and the solution should be between 0 and 2π 29 is a little too big...

OpenStudy (anonymous):

In degrees \[ 0 \le x \le 360 \]

OpenStudy (anonymous):

radians

OpenStudy (anonymous):

here's the answers i see in my book; x=.508, 1.06, 3.65, 4.20

OpenStudy (anonymous):

You should get 4 solutions

OpenStudy (anonymous):

how am i suppose to get 4 solutions? am i suppose to graph it and see when it intersects the 0?

OpenStudy (anonymous):

Here is the graph

OpenStudy (anonymous):

\[ \sin(u) = \sin (\alpha) \implies u = \alpha + 2 k \pi \text { or } u =\pi- \alpha + 2 k \pi \]

OpenStudy (anonymous):

\[\sin^{-1} 0.85=1.016 ,2x=1.016,x=0.508\] x=0.508 radians

OpenStudy (anonymous):

oh hey thanks, the graph helped me. i graphed it on my calculator and found where it intersects. y1= sin2x y2= .85 is that how you did it? anyways thanks for the help :)